OpenStudy (anonymous):

which of the following is the equation of a line that has slope -3 and passes through (4,-5)

OpenStudy (anonymous):

3x+y-7=0

OpenStudy (anonymous):

-3x + 7 = y

OpenStudy (amistre64):

y = -3(x-4)-5 y = -3x +12 -5 y = -3x +7

OpenStudy (anonymous):

y+5=-3(x-4) y+5=-3x+12 3x+y=7

OpenStudy (amistre64):

the slope point form is just as valid an equation for the line as the ret too ;)

OpenStudy (anonymous):

ret?

OpenStudy (amistre64):

shorthand for rest lol

OpenStudy (anonymous):

yeah, you are going to use the point-slope form.

OpenStudy (amistre64):

nah, ill just use yours: y+5=-3(x-4)

OpenStudy (anonymous):

thanks.